# Vistoria: A Multimodal System to Support Fictional Story Writing through Instrumental Text–Image Co-Editing

### v2

@@ Abstract.
Humans think visually—we remember in images, dream in pictures, and use visual metaphors to communicate. Yet, most creative writing tools remain text-centric, limiting how authors planing and translating ideas. We present Vistoria, a system for synchronized text–image co-editing in fictional story writing that treats visuals and text as co-equal narrative materials. A formative Wizard-of-Oz co-design study with 10 story writers revealed how sketches, images, and annotations serve as essential instruments for ideation and organization. Drawing on theories of Instrumental Interaction and Structural Mapping, Vistoria introduces multimodal operations-lasso, collage, filters, and perspective shifts that enable seamless narrative exploration across modalities. A controlled study with 12 participants shows that co-editing enhances expressiveness, immersion, and collaboration, enabling writers to explore divergent directions, embrace serendipitous randomness, and trace evolving storylines. While multimodality increased cognitive demand, participants reported stronger senses of authorship and agency. These findings demonstrate how multimodal co-editing expands creative potential by balancing abstraction and concreteness in narrative development.
Multimodality, Creativity Support, Storytelling, Creative Writing, Instrumental Interation, Structure Mapping, Direct Manipulation
journalyear: 2026copyright: acmlicensedconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Creant el demà junts, Barcelonabooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Creant el demà junts, Barcelona
Figure 1. Vistoria supports a cyclic workflow in which multimodal artifacts and ideas co-evolve. (A) Instrument Interaction: a unified set of tools—Collage, Lasso, Filter, and Perspective Shift—enables coupled editing of image and text to produce image–text “cards.” (B) Artifacts → Idea: the resulting visual and textual artifacts stimulate new concepts and story directions. (C) Cognitive Support: a structural mapping view aligns image and text variations, making alternatives comparable and keeping narrative structure consistent. (D) Ideas → Artifacts: emerging ideas are materialized back into new cards, closing the loop and driving iterative ideation, exploration, and integration.
@@ 1. Introduction
Human cognition is deeply rooted in visual processing (Rezaee, 2011; Huang et al., 2024; Kanellopoulou et al., 2019). We recall experiences as spatial scenes, build mental models through imagery, and rely on techniques such as memory palaces to organize abstract information. Everyday reasoning often involves manipulating internal visual representations—planning a route, rearranging objects in mind, or picturing alternative outcomes. Language use is similarly tied to imagery—comprehension frequently evokes mental pictures, and abstract ideas are often conveyed through spatial metaphors such as path, framework, or perspective (Meng et al., 2023). This aligns with Dual Coding Theory, which argues that cognition integrates both verbal and nonverbal channels, with distinct representational capacities, to support imagination, reasoning, and communication (Clark and Paivio, 1991).
Given this inherent multimodality of thought, a critical question arises: why do our most sophisticated creative practices, such as narrative writing, remain confined to unimodal tools? Fictional story writers often describe richly visualized scenes and manipulate complex spatial relationships in their imagination, yet must translate these processes through the narrow channel of sequential text (Mukramah et al., 2023; Listyani, 2019; Rao et al., 2024; Wang et al., 2025c). This mismatch between the multimodal nature of creative cognition and the unimodal design of creative tools constrains expressive potential (Zhao et al., 2025; Lee et al., 2024).
In fictional story writing, the most demanding phases are planning and translating (Gero et al., 2022; Flower and Hayes, 1981). Authors must organize imagined scenes, characters, and events into coherent structures, then render these abstractions into linear prose (Doyle, 1998; Bal and Veltkamp, 2013; Chung et al., 2025; Dang et al., 2023). Visuals can scaffold these processes by stimulating ideas, supporting organization, and enriching narrative detail (Mukramah et al., 2023; Listyani, 2019; Ali and Parikh, 2021). However, traditional visual aids have been limited in scope and scalability, offering only passive inspiration rather than dynamic, interactive support for the iterative writing process (Kreminski et al., 2024; Choi et al., 2024; Yan et al., 2023; Mukramah et al., 2023; Listyani, 2019). Large language models (LLMs) have unlocked unprecedented generative capabilities, enabling rapid creation of diverse narrative elements through combined text and visual imagery at previously unattainable scales (Fu et al., 2024). LLM-powered systems have explored visual integration through character construction tools (Qin et al., 2024), visual analytics for branching storylines (Mishra et al., 2025; Wang et al., 2025a), script visualization (Rao et al., 2024), diagram-prose synchronization (Masson et al., 2024b), and text-editing metaphors from image manipulation (Masson et al., 2024a).
However, the use of visuals remains peripheral across these approaches. They are often treated as prompts, overlays, or loosely coupled representations with unidirectional interaction. Our formative Wizard-of-Oz (WOz) co-design study with 10 writers revealed that sketches, annotations, and image manipulations are essential instruments for externalizing mental images and organizing narrative ideas. What is missing from existing LLM-powered creativity tools for fictional story writing is a unified multimodal substrate that directly supports iterative oscillation between planning and translating, where text and visuals serve as co-equal, manipulable materials in a bidirectional co-editing loop.
To bridge this gap, we developed Vistoria, a system that transforms fictional story writing from a unimodal text process into a multimodal co-editing experience where synchronized text-image manipulations through polymorphic operations—lasso, collage, perspective shifts, and filters. The design enables writers to fluidly navigate between abstract conceptualization and concrete visualization while creating a unified multimodal substrate for narrative development. These operations enable writers to reorganize narrative elements, explore alternative viewpoints, and refine story details through a bidirectional loop where text and visuals continuously inform one another. The design of Vistoria was grounded in two key theoretical frameworks. Instrumental Interaction guides consistent operations across modalities to reduce switching costs and sustain flow (Beaudouin-Lafon, 2000; Shi et al., 2025; Riche et al., 2025), while Structural Mapping (Gentner, 1983; Norman, 2013) explains how verbal and visual representations can be aligned so edits in one modality inform the other.
We conducted a controlled study with 12 participants to evaluate Vistoria’s effectiveness in supporting multimodal narrative creation. The results indicate that participants successfully integrated Vistoria’s synchronized text-image co-editing into their creative workflows, employing four distinct multimodal strategies to enhance storytelling. Participants demonstrated enhanced creative expression by using images as catalysts for richer descriptions, leveraging visual editing operations to drive narrative pivots, and employing multimodal elements to externalize and interconnect story branches. The cross-modal dialogue facilitated creativity through divergent exploration of multiple narrative directions, serendipitous discoveries from AI-generated content, and culture-driven associations that activated personal memories and references. While multimodality increased cognitive load, study participants reported heightened authorship, agency, and immersion, characterizing the system as a collaborative partner that expanded their imagination. Ultimately, creativity emerged from the dynamic interplay between modalities, where text and imagery continuously shaped one another.
@@
  • A WOz co-design study with 10 writers examined the practices and needs of using multimodal elements (text, sketches, and visual imagery) to externalize ideas and develop narratives in the planning and translating phase of fictional story writing;
  • Vistoria, a multimodal co-editing system that unifies text and visual images through polymorphic operations and a bidirectional editing loop to support planning and translating in fictional story writing;
  • A controlled usability study with 12 participants demonstrates the effectiveness of Vistoria, showing that multimodal co-editing enhances expressiveness, coherence, and authorial agency in fictional story writing.
@@ 2.1. Using Visuals to Support the Cognitive Process of Fictional Story Writing
Fictional story writing is distinct from argumentative or expository genres in its emphasis on imagination, world-building, and character development (Doyle, 1998). Writers must invent narrative worlds and characters while ensuring coherence, which poses unique cognitive challenges: abstract, non-verbal mental images must be transformed into structured narrative elements and then into text (Bal and Veltkamp, 2013).
@@
The Cognitive Process Model of Writing (Gero et al., 2022; Flower and Hayes, 1981) frames writing as recursive processes of planning, translating, and reviewing. In fictional story writing, the transition from planning to translating is especially demanding, as writers move from imaginative constructs to linear verbal representation, imposing a high cognitive load. However, visual representations can scaffold this process by externalizing abstract ideas. Research shows that picture prompts improve writing coherence (Mukramah et al., 2023), and visual images stimulate creativity in narrative writing (Listyani, 2019). In practice, sketches, maps, and diagrams help structure plot, setting, and character relationships during planning, and serve as cognitive anchors during translating to support coherence and consistency (Zhao et al., 2025). Dual Coding Theory (Clark and Paivio, 1991) explains these benefits: verbal and imagery systems function separately but also interact, creating richer memory traces when information is encoded in both modalities. In fictional writing, visual representations of narrative elements complement verbal planning, making abstract concepts more concrete and retrievable. When writers encounter difficulties in translation, visual anchors provide alternative access to imaginative content, reducing cognitive load and enabling more fluid expression (Barsalou, 2008).
@@
Building on this foundation, a variety of creativity support tools have incorporated visuals into different phases of writing. Planning-focused systems such as CCI, Sketchar, and CharacterMeet assist authors in character and world development, often through image-guided backgrounds or conversational refinement of characters (Park et al., 2024; Long et al., 2024; Qin et al., 2024). Translation-focused tools like ScriptViz and Script2Screen aim to align textual composition with visual referents, either by retrieving reference visuals from movie databases (Rao et al., 2024) or by synchronizing scriptwriting with audiovisual scene creation (Wang et al., 2025c). Complexity management systems, for example, WhatIF (Mishra et al., 2025), ClueCart (Wang et al., 2025a), and PlotMap (Wang et al., 2024), help writers maintain structural coherence by visualizing branched narratives, organizing narrative clues hierarchically, or integrating spatial layouts with textual plot structures.
Despite these advances, visuals in prior systems remain peripheral: they serve primarily as references, scaffolds, or analytic overlays, rather than as co-equal, manipulable materials. This treatment constrains multimodal creative expression by limiting writers to either text-first workflows or loosely coupled visual aids (Zhao et al., 2025). To bridge this gap, our work introduces a system that treats visuals and text as fully integrated, co-editable artifacts, directly supporting the planning and translating phases of fictional story writing so that imagination can be externalized and refined seamlessly across both modalities.
@@ 2.2. LLM-powered Multimodality Tools for Creativity in Content Creation
Recent multimodal creativity tools move beyond linear prompting by enabling direct manipulation of creative elements, helping creators express intentions that language alone cannot capture (Riche et al., 2025; Masson et al., 2024a). Powered by advanced LLMs, these systems address fundamental barriers through three complementary mechanisms. First, externalizing creative structures. Tools like AI-Instruments (Riche et al., 2025) and Brickify (Shi et al., 2025) turn abstract intentions into manipulable interface objects or reusable design tokens, making spatial, stylistic, and compositional relations tangible. Second, reducing cognitive load with multimodal input: DrawTalking (Rosenberg et al., 2024) and Code Shaping (Yen et al., 2025) distribute expression across sketch, voice, or annotations, lowering the reliance on precise verbal specification. Third, supporting iterative refinement. Inkspire (Lin et al., 2025) and AIdeation (Wang et al., 2025b) accelerate variation and exploration, enabling rapid cycles of sketch-to-output or recombination of references. However, most current systems remain fragmented: they either support isolated stages (e.g., ideation, editing) or couple modalities only loosely, often within a one-shot prompting paradigm that underutilizes LLMs’ potential for sustained dialogue and refinement (Dang et al., 2023).
Seeking tighter coupling for fictional story writing, recent systems push integration of multimodal interaction in different ways (Chung et al., 2025; Chung and Kreminski, 2024). WorldSmith grounds worldbuilding in text–sketch edits and region-based filling to support users iteratively visualize and modify elements of their fictional world (Dang et al., 2023). XCreation (Yan et al., 2023) supports cross-modal storybook creation by integrating an interpretable entity-relation graph, improving the usability of the underlying generative structures. Toyteller (Chung et al., 2025) maps symbolic motions to character actions. Visual Writing allows writers co-edit stories by directly manipulating visual representations of entities, actions, and events alongside text, enabling easier navigation, editing, and exploration of narrative variations (Masson et al., 2024b). These systems highlight the transformative potential of LLMs to couple narrative and visual modalities, yet they still lack a unified, synchronized substrate that supports fictional story writing as a continuous, multimodal workflow. Our work advances this direction by introducing such a substrate, where LLM-generated texts and visuals act as co-editable inspirational materials for iterative, coherent story development.
2.3. Instrumental Interaction and Structural Mapping Theory

### v3

# Vistoria: A Multimodal System to Support Fictional Story Writing through Instrumental Image-Text Co-Editing

@@ Abstract.
Humans think visually—we remember in images, dream in pictures, and use visual metaphors to communicate. Yet, most creative writing tools remain text-centric, limiting how writers plan and translate ideas. We present Vistoria, a system for synchronized image-text co-editing in fictional story writing. A formative Wizard-of-Oz co-design study with 10 story writers revealed how sketches, images, and text serve as essential elements for ideation and organization. Drawing on theories of Instrumental Interaction, Vistoria introduces instrumental operations—Lasso, Collage, Perspective Shift, and Filter that enable seamless narrative exploration across modalities. A controlled study with 12 participants shows that co-editing enhances expressiveness, immersion, and collaboration, opening space for writers to follow divergent story directions and craft more vivid, detailed narratives. While multimodality increased cognitive demand, participants reported stronger senses of ownership and agency. These findings demonstrate how multimodal co-editing expands creative potential by balancing abstraction and concreteness in narrative development.
Multimodality, Creativity Support, Storytelling, Creative Writing, Instrumental Interaction
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790400isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Interaction paradigms
Figure 1. Vistoria supports a cyclic workflow in which multimodal artifacts and ideas co-evolve. (A) Instrumental operations enable image–text co-editing. (B) Resulting image–text artifacts prompt new story directions. (C) Image–text alignment supports idea formation by coordinating verbal and non-verbal processing. (D) Emerging ideas are externalized into new artifacts, closing the loop for iterative ideation and integration.
The figure illustrates Vistoria’s overall cyclic workflow for multimodal story writing. It shows how writers interact with image–text instruments to edit visual and textual artifacts, how these artifacts stimulate new story ideas, and how ideas are materialized back into new image–text cards. The cycle highlights continuous co-evolution between artifacts and ideas, emphasizing synchronized visual and verbal processing during iterative narrative development.
@@ 1. Introduction
Human thinking involves multimodal processing. Visual processes play a central role in cognition: we recall experiences as spatial scenes, form mental models through imagery, and use visual structure to organize and interpret information  (Rezaee and Shoar, 2011; Huang et al., 2024; Kanellopoulou et al., 2019). Language is similarly entwined with imagery. Particularly, text comprehension often evokes mental pictures, and abstract ideas are commonly articulated through spatial metaphors such as path, framework, or perspective (Zhao et al., 2025; Clark and Paivio, 1991; Gärdenfors, 1996). Dual Coding Theory frames this coupling between imagery and language, positing that humans draw on both verbal and nonverbal channels to represent and support specific reasoning and communicative processes (Clark and Paivio, 1991).
Story writing is particularly multimodal in nature. During the planning phase, experienced writers often use both imagery and language to construct the story world. They visualize spatial layouts, character interactions, and scene dynamics, while using textual notes to label, sequence, and reason about narrative structure (Doyle, 1998; Bal and Veltkamp, 2013; Chung et al., 2025; Dang et al., 2023). In the translating phase, visual details serve as an anchor that shapes how writers organize narrative detail and emotional tone, while texts linearize these visualized ideas into descriptions, dialogue, and narrative perspective that readers can follow (Mukramah et al., 2023; Listyani, 2019). Crucially, nonverbal and verbal channels do not operate in isolation: writers use imagery to trigger new wording, and emerging text in turn elicits further mental images (Zhao et al., 2025; Lee et al., 2024). Therefore, there is a great opportunity for tools that support story writing to match this multimodal complexity by accommodating the continuous interplay between visual and textual thinking within the same workflow. Yet, current writing tools remain overwhelmingly text-centric, treating linear text as the primary or sole medium of expression (Lee et al., 2024; Gero et al., 2023). Although recent systems powered by Large Language Models (LLMs) incorporate visual elements through image generation or retrieval, these visuals remain peripheral. They function mainly as prompts (Qin et al., 2024), static references (Rao et al., 2024), or organizational diagrams (Mishra et al., 2025; Wang et al., 2025b), rather than as tightly integrated, manipulable representations along with text. This requires writers to translate visual ideas back into text, increasing cognitive load (Brown and Kim, 2021; Zhao et al., 2025; Yan et al., 2023; Singh et al., 2023). Such translation overhead not only hinders cross-modal collaboration but also restricts the creative synthesis of semantically distant concepts (Singh et al., 2023) that would emerge more naturally in a truly integrated multimodal system.
To examine this gap, we conducted a formative Wizard-of-Oz (WoZ) co-design study with 10 experienced writers to inform the design of a unified image-text multimodal system that supports the planning and translation phases of fictional story writing. We found that text and images play distinct yet complementary roles in the writing process. Writers expressed a need to directly manipulate text and images for fine-grained editing and alignment, valuing the ability to move fluidly between them.
Based on these formative results, we developed Vistoria, a system for experienced, LLM-literate writers that reimagines fictional storytelling as a multimodal co-editing experience, integrating text, images, and sketches beyond traditional text-centric workflows. The design of Vistoria draws on the principle of Instrumental Interaction, designing a set of instrumental operations (lasso, collage, perspective shift, and filter). These functions can be applied to either text or images, where a single action simultaneously affects both images and text, minimizing switching costs and preserving creative flow (Beaudouin-Lafon, 2000; Shi et al., 2025; Riche et al., 2025). Based on Dual Coding Theory, Vistoria enables the alignment of text and visual representations, ensuring that edits in one modality are appropriately reflected in the other.
We conducted a controlled study with 12 participants to examine how multimodal image–text co-editing supports fictional story writing, with a focus on evaluating the system’s usability and understanding its creative support. Overall, the study showed that Vistoria enhances expressiveness, immersion, and exploration, enabling participants to have more divergent ideas and write detailed narratives. Participants used the instrumental operations to refine ideas at multiple scales and explore alternative directions. While this workflow increased mental and physical workload, it also supported writers’ senses of agency and ownership, as they maintained greater operational control over narrative development.
@@
  • A WoZ co-design study with 10 writers examined the practices and needs of using multimodal elements to externalize ideas and develop narratives in the planning and translating phase of fictional story writing;
  • Vistoria, a multimodal co-editing system designed for experienced, LLM-literate writers that unifies text and visual images through instrumental operations and a synchronized editing loop to support fictional story writing;
  • A controlled usability study with 12 participants demonstrates the potential of Vistoria, suggesting that multimodal co-editing can enhance expressiveness, idea generation, and narrative development in fictional story writing.
@@ 2.1. Using Visuals to Support the Cognitive Process of Fictional Story Writing
Fictional story writing is distinct from argumentative or expository genres in its emphasis on imagination, world-building, and character development (Doyle, 1998). Writers must invent narrative worlds and characters while ensuring coherence, which poses unique cognitive challenges: abstract, nonverbal mental images must be transformed into structured narrative elements and then into text (Bal and Veltkamp, 2013).
@@
The Cognitive Process Model of Writing (Gero et al., 2022; Flower and Hayes, 1981) frames writing as recursive processes of planning, translating, and reviewing. In fictional story writing, the phases from planning to translation are especially demanding, as writers move from imaginative constructs to linear verbal representation, imposing a high cognitive load due to the need for simultaneous translation and structural organization. However, visual representations can scaffold this process by externalizing abstract ideas. Research shows that picture prompts improve writing coherence (Mukramah et al., 2023), and visual images stimulate creativity in narrative writing (Listyani, 2019; Han et al., 2024). In practice, sketches, maps, and diagrams externalize plot, setting, and character relationships that writers actively manipulate during planning and revision, while also serving as cognitive anchors during translation to maintain coherence and consistency. (Zhao et al., 2025). Dual Coding Theory (Clark and Paivio, 1991) explains these benefits: verbal and nonverbal systems function separately but also interact, creating richer memory traces when information is encoded in both modalities. In fictional writing, visual representations of narrative elements complement verbal planning, making abstract concepts more concrete and retrievable. When writers encounter difficulties in translation, visual anchors provide alternative access to imaginative content, reducing cognitive load and enabling more fluid expression (Barsalou, 2008). These visual structures are not merely supportive; they function as alternative representational spaces in which writers perform cognitive operations that parallel textual editing and support non-linear narrative leaps (Yan et al., 2023; Singh et al., 2023).
@@
However, existing creativity-support systems largely leverage visuals in limited ways, focusing on inspiration, reference, or structural overview rather than enabling writers to directly manipulate visual narrative elements and align with text editing. Planning-focused systems such as CCI, Sketchar, and CharacterMeet assist authors in character and world development, through image-guided backgrounds or conversational refinement of characters (Park et al., 2024; Ling et al., 2024; Qin et al., 2024). Translation-focused tools like ScriptViz and Script2Screen aim to align textual composition with visual referents, either by retrieving reference visuals from movie databases (Rao et al., 2024) or by synchronizing scriptwriting with audiovisual scene creation (Wang et al., 2025c). Complexity management systems, for example, WhatIF (Mishra et al., 2025), ClueCart (Wang et al., 2025b), and PlotMap (Wang et al., 2024), help writers maintain structural coherence by visualizing branched narratives, organizing narrative clues hierarchically, or integrating spatial layouts with textual plot structures.
In these systems, writers may look at images to spark ideas, but operations such as cutting, re-ordering, or reframing narratives still have to be performed only in the verbal channel. Because the underlying story state is effectively defined only through text, visuals cannot serve as core writing operations such as restructuring events, adjusting focalization, or reorganizing character relationships (Flower and Hayes, 1981; Zhao et al., 2025). This separation requires writers to repeatedly convert visually grounded ideas back into verbal form for any narrative change to take effect, thereby increasing cognitive load and undermining many of the well-established benefits of external representations such as diagrams, sketches, and other forms of external cognition (Kirsh, 2010; Singh et al., 2023). As a result, images remain outside the recursive planning–translating loop. To address this gap, our work treats visual and text representations as synchronized and co-editing materials, allowing writers to manipulate narrative elements across image and text through the same set of operations and thereby more tightly aligning verbal and nonverbal with the cognitive processes of fictional story writing.
@@ 2.2. LLM-powered Multimodality Tools for Creativity in Content Creation
Recent multimodal creativity tools move beyond linear prompting by enabling direct manipulation of creative elements, helping creators express intentions that language alone cannot capture (Riche et al., 2025; Masson et al., 2025a). Powered by LLMs, these systems address fundamental barriers through three complementary mechanisms. First, they externalize creative structures and support better intention expression. Tools like AI-Instruments (Riche et al., 2025) and Brickify (Shi et al., 2025) transform abstract intentions into manipulable interface objects or reusable visual tokens, rendering otherwise ineffable ideas as visible, persistent, and operable elements. Second, in recent multimodal systems, sketches always act as a nonverbal, spatially grounded modality that conveys structure, hierarchy, and relations far more efficiently than language. DrawTalking (Rosenberg et al., 2024) combines freehand sketching with spoken narration, enabling natural intention communication. Code Shaping (Yen et al., 2025) allows developers to make sketched annotations directly on the code editor to support fuzzy, incremental expression of intent. Third, supporting iterative refinement, Inkspire (Lin et al., 2025) and AIdeation (Wang et al., 2025a) accelerate variation and exploration, enabling rapid cycles of sketch-to-output or recombination of references. Similarly, the AI Drawing Partner (Davis and Rafner, 2025) combines sketch recognition, text-to-image generation, and direct feedback mechanisms to enable rapid cycles of creative exploration while automatically quantifying the interaction dynamics between human and AI collaborators.
Seeking tighter coupling for fictional story writing, recent systems push the integration of multimodal interaction in different ways (Chung et al., 2025; Chung and Kreminski, 2024). These systems are designed in response to growing evidence that text-only, model-driven workflows cause LLM-assisted stories to converge toward similar narrative structures, limiting exploration, reducing originality, and diminishing writers’ expressive control (Chakrabarty et al., 2024a; Doshi and Hauser, 2024; Li et al., 2024). WorldSmith supports layered edits and hierarchical compositions through sketches, making it easier to grow a world piece by piece instead of through single and monolithic prompts (Dang et al., 2023). XCreation (Yan et al., 2023) supports cross-modal storybook creation by integrating an interpretable entity-relation graph, improving the usability of the underlying generative structures. Toyteller (Chung et al., 2025) maps symbolic motions to character actions, letting users express rich social and emotional interactions that are often hard to write down explicitly only using text. Visual Writing defines an approach where writers edit stories by manipulating visual representations to make the underlying narrative structure more comprehensible and easier to work with than linear text alone (Masson et al., 2025b).
This line of multimodal research demonstrates that combining language with gestures, sketches, and direct manipulation can offload cognitive work from linear prompting and give creators more expressive, situated channels for specifying and revising intent. Building on this line of work, our system introduces a canvas-based interface that integrates images, sketches, and text to support the externalization of mental imagery and the articulation of narrative intent.
2.3. Instrumental Interaction